Q. The osmotic pressure of a solution of a synthetic polyisobutylene in benzene was determined at 25°C. A sample containing 0.20 g of solute/100 cm3 of solution developed a rise of 2.4 mm at osmotic equilibrium. The density of the solution was 0.88 g/cm3. What is the molecular weight of the polyisobutylene?

Grade:

1 Answers

Naveen Kumar
9 years ago
Osmotic pressure= C*R*T
Lat molecular weight ofpolyisobutylene=M
Hence concentration=(0.2/M)*10=Molarity ofpolyisobutylene
Now we know that 760mm of Hg=1atm,
So 2.4 mm of Hg=1*2.4/760atm= 0.0032
Hence, 0.0032={(0.2/M)*10}*0.082*298 and calculaate value of M